Design the buck and buck-boost converters for following specifications: Vin = 24 V, Switching frequency = 25 kHz, Vo = 10 V, Po= 48 W. Inductor current ripple is 10 % of load current and capacitor voltage ripple is 1% of output voltage. Assume necessary data if required.


1
Expert's answer
2021-04-26T02:03:57-0400

"V_o=V_s(\frac{D}{1-D})"

"10=24(\frac{D}{1-D}) \implies D=0.29411"

"L_{min}=\frac{(1-D)^2R}{2f}=\frac{(1-0.29411)^2 \times 24^2/48}{2 \times 25}=0.119587"

Required Inductance "1.51 \times L_{min}=1.51 \times 0.119587=0.18057 \mu H"

"C=\frac{D}{Rf(V)}=\frac{0.29411}{24^2/48(38/48)}=0.0309589 \mu F"